OpenStudy (anonymous):

Write as an algebraic expression in x: cos(arcsin(5/13))

OpenStudy (anonymous):

draw a triangle with opposite side 5 and hypotenuse 13 find the other leg via pythagoras that is all

OpenStudy (anonymous):

the other leg is 13.928?

OpenStudy (anonymous):



rvc (rvc):

a^2+5^2=13^2

OpenStudy (anonymous):

yeah, so b=13.928?

OpenStudy (anonymous):

Then what?

rvc (rvc):

what is the value of b^2 ?

OpenStudy (anonymous):

194

OpenStudy (anonymous):

oh no

rvc (rvc):

169-25=?

OpenStudy (anonymous):

the other leg is less than the hypotenuse

rvc (rvc):

b^2+ 25 = 169 b^2 = 169-25

OpenStudy (anonymous):

144

rvc (rvc):

so b=?

OpenStudy (anonymous):

12

OpenStudy (anonymous):

then what?

rvc (rvc):

what next?

OpenStudy (anonymous):

I don't know. That's why I'm asking.

rvc (rvc):

@satellite73 could u please continue

Nnesha (nnesha):

\[\rm sin \rm \theta = \frac{ opposite }{ hypotenuse }~~~~ \cos \theta = \frac{ adjacent }{ hypotenuse } ~~\tan \theta = \frac{ opposite }{ adjacent }\]  cos = ?

OpenStudy (anonymous):

12/13

OpenStudy (anonymous):

@Nnesha

Nnesha (nnesha):

:-) yes

OpenStudy (anonymous):

But what do I do with that?

Nnesha (nnesha):

put some jam on it and then eat with c (chocolate ) :-) that's it !!! :-)

OpenStudy (anonymous):

but the answer in my calculator is different from my answer (again)!

Nnesha (nnesha):

;-; what's your answer :?

rvc (rvc):

cos(cos(pi/2-12/13))

OpenStudy (anonymous):

what?

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @AEB047 but the answer in my calculator is different from my answer (again)! \(\color{blue}{\text{End of Quote}}\) what u got ?

rvc (rvc):

sin(x)=cos(pi/2-x)

OpenStudy (anonymous):

.9230769231 on calculator .999870225 by me

Nnesha (nnesha):

0.92307 =12/13

rvc (rvc):

:)

OpenStudy (anonymous):

It's the same if I put in (cos(arcsin(5/13))!

rvc (rvc):

All the best

Nnesha (nnesha):

http://prntscr.com/76usl3

Nnesha (nnesha):

why do you need in decimal form ? keep it as fraction

OpenStudy (anonymous):

Oh, so I put in 5/13 into the cos(arcsin but 12/13 into my equation.

rvc (rvc):

what is ur question???

Nnesha (nnesha):

cos(arcsin(5/13)) = 12/13

Nnesha (nnesha):

arcsin is a 5/13

OpenStudy (anonymous):

Ohhhhh!!! I think I get it now! Thanks!

rvc (rvc):

:)

OpenStudy (anonymous):

My answer is just 12/13!

Nnesha (nnesha):

glad :-)

Nnesha (nnesha):

yeah